# Modeling Short-Range Microwave Networks to Scale Superconducting Quantum Computation

### v3

@@ Abstract
A core challenge for superconducting quantum computers is to scale up the number of qubits in each processor without increasing noise or cross-talk. Distributed quantum computing across small qubit arrays, known as chiplets, can address these challenges in a scalable manner. We propose a chiplet architecture over microwave links with potential to exceed monolithic performance on near-term hardware. Our methods of modeling and evaluating the chiplet architecture bridges the physical and network layers in these processors. We find evidence that distributing computation across chiplets may reduce the overall error rates associated with moving data across the device, despite higher error figures for transfers across links. Preliminary analyses suggest that latency is not substantially impacted, and that at least some applications and architectures may avoid bottlenecks around chiplet boundaries. In the long-term, short-range networks may underlie quantum computers just as local area networks underlie classical datacenters and supercomputers today.
I Introduction
@@
Quantum computing is at the forefront of emerging technologies, owing to its promise to solve some algorithms dramatically faster than classical counterparts. Quantum computers (QCs) evaluate quantum circuits in a manner similar to classical computers assessing classical circuits. However, quantum information’s ability to leverage superposition, interference, and entanglement is anticipated to provide QCs significant advantage in tasks ranging from security protocols [1] and optimization [2], to simulation of quantum dynamics [3]. To harness the power of quantum computation, quantum bits, or qubits, are needed which are the core individual element for information representation and storage. Useful QCs require many qubits that are interconnected to each other in a low-noise and fault-tolerant manner. Many different architectures have been used to encode qubits, such as superconducting circuits (SCs) [4, 5, 6], trapped ions [7], neutral atoms [8], quantum dots [9], and photons [10, 11]. To date, SC quantum computers have been one of the most promising platforms, with the largest quantum computers in terms of interconnected qubits demonstrated with these circuits. Recent demonstrations of quantum advantage have shown that interference with just over 50 qubits can be harnessed for computational speedups [4, 6, 10]. Additionally, early demonstrations of logical qubits are emerging [12, 13]. However, more and cleaner qubits will be required to exploit quantum computing for practical applications.
@@
State of art QCs, sometimes called Noisy Intermediate Scale Quantum (NISQ) devices, are error prone are reaching a thousand-qubit capacities [14]. The qubits on these QCs are not error corrected and often have limited connectivity. NISQ qubits are characterized by their sensitivity to interactions with their environment and each other, and the gates and measurements required for meaningful computation are implemented with imperfect, error-prone operations. To achieve the realm of usefulness for quantum devices, many more qubits, approximately one million or more, and lower noise figures are essential [15]. Concerningly, as noted in [15], “quantum computer designers face a trade-off between trying to optimize for quantum computers with many physical qubits and quantum computers with very low gate error rate.”
@@
Classical computing saw similar challenges associated with scaling the system-on-chip (SoC) model [27]. Integrated circuit (IC) yield loss is a result of manufacturing defects caused randomly by environmental particulates or by systematic sources such as unavoidable variation or errors during fabrication [28]. Because a larger chip has a higher probability of containing a defect that renders the entire system unusable, fabricating fewer, large SoCs on a wafer results in lower yields than ICs with a smaller wafer footprint. This observation has led to the increase in popularity of chiplet-based architectures for improved cost and flexibility in system design and use [29]. Just as pitfalls were eventually observed for monolithic classical computing devices, it is anticipated that continuing to pack qubits on single chip will push the capabilities of manufacturing, cooling, and control technologies to their limit. Thus, to develop QCs containing the millions of qubits required fault-tolerance, we are motivated to pursue a modular approach based on multiple cores working in sync to accomplish computational tasks in a distributed manner.
@@
I.1 Main Contributions & Where This Work Fits
A key contribution of this work is to bridge the gap from understanding microwave links at the physical layer to using them in distributed quantum computing applications. Rather than encapsulate network overheads as abstract cost functions, we analyze how noise and costs in both cross-chiplet links and local processing impact the eventual fidelity and feasibility of computations. Our models start and extrapolate from observations of real hardware and experimental physics models. We propose practical, concrete schemes to detect and reduce errors on near-term links and processors. Broad themes in potential applications are evaluated for feasibility on devices with limited connectivity.
@@ I.1 Main Contributions & Where This Work Fits
The structure of this paper is as follows: In Section II, we review some related literature including analogous approaches to similar questions. Section III is about modeling quantum links and local gates at the physical layer, including a comparison of Hamiltonians between optical and microwave links, a review of hardware challenges for local processing, and quantum channel models of relevant operations. We present a quantum channel model that captures realistic noise on both microwave links and local SWAP operations, setting up later comparions. In Section IV, we compare monolithic with chiplet architectures at the network layer, incorporating physical information to derive expected fidelities of qubit transfers in addition to topological metrics. We also derive a detailed channel model that includes both links and local swaps, which we use to propose a simple error detection scheme for the regime of high-quality local processing with lossy links. Finally, we conclude with some open problems and areas for potential follow-up research.
II Background
@@
II.1 Two major regimes in quantum networking
@@
II.2 Ongoing work in short-range link hardware
@@
In Section III, we describe the physics and characteristics of these links, comparing them with matter-optical transduction.
II.3 Algorithms and applications
@@
III Physical components of a distributed network
III.1 On-chip hardware
@@ III.2 Link hardware
III.3 Modeling link-based and local SWAP operations
@@
We may alternatively regard a SWAP operation or link as a pair of single-qubit channels, moving a qubit state from the $i$th to $j$th physical qubit. Though this picture is less complete, it vastly simplifies the analysis of multiple moves as in Section IV, characterizing a single-qubit channel via tomography rather than a channel of dimension exponential in the length of the SWAP chain. Empirically, we observed in the experiments for Section IV.2.2 that a long chain model composed from single qubit SWAP tomographies more accurately recovers that of a long chain than does a reduced two-qubit model. We believe this discrepancy reflects cross-talk in nearby measurements or complications in readout error mitigation. The single qubit channel probably characterizes the SWAP more accurately by incorporating less readout noise. Here we again determine an average Choi matrix based on data from the 27-qubit $ibmq\_montreal$:
@@
IV Main Results: modeling distributed architectures and comparing to monolithic
@@
IV.1 Connectivities and graph metrics
@@
    As a monolithic comparison point, we consider the standard two-dimensional grid, in which each qubit not on the boundary has four neighbors.
@@
    We split the heavy-hex layout into chiplets by replacing some connections with links. This has little effect on topology but may change the error tradeoffs as described in Subsection IV.2. In particular, we partition the chiplet into 80-qubit blocks, each with up to 16 links for a qubit-to-link ratio of at least 10. A smaller version of this layout is shown in Figure 2(c). Unlike the grid case, here we can maintain a reasonable qubit-to-link ratio while keeping the heavy-hex structure across the entire device consistent. We expect topological properties similar to the monolithic heavy-hex but impose a rectangular overall structure, which might be less square than its monolithic counterpart due to the minimum block size.
@@
Figure 4: Comparison of QC topology spectral gap vs. number of qubits for proposed architectures. The spectral gap as in Definition IV.2 quantifies the rate at which random walks on the graph converge toward a uniform distribution on nodes. Graphs with few bottlenecks tend to have higher spectral gaps.
@@
Although topological metrics appear to favor the monolithic grid over other planar layouts, this does not necessarily imply superiority as a quantum computing architecture. As noted in [20], heavy-hex and related layouts achieve lower coordination numbers, taking pressure off of other considerations and potentially enabling technique for higher fidelity.
@@
IV.1.1 Bandwidth Considerations for Applications
@@
In entanglement forging [61] , an observable on a whole system is rewritten as a linear (not necessarily convex) combination of observables on two subsystems via Schmidt decomposition. Entanglement forging simulates the expectation of an observable $\braket{\mathcal{O}}_{\ket{\psi}}$ on an $n$-qubit system with operator norm error $\epsilon$ via repeated uses of an $n/2$-qubit system. The distributed scheme would use two $n/2$-qubit systems and a number of link uses proportional to the entanglement needed between subsystems. In principle, if each bidirectional link use exchanges up to two qubits of entanglement, the number of link uses is logarithmic in the number of non-zero Schmidt coefficients. For comparability, one should allow a chiplet mapping to approximate the true state within precision $\epsilon$ as well. If we post-select on successful transfers via the two-use code discussed in Section IV.4, then we may convert link uses and noise into a failure probability. The chance that each double-transfer fails is $1/\eta$, where $\eta$ is an exponentiated amplitude damping parameter. In this formulation, we can directly compare the number of retrials expected using $\frac{1}{2}\log_{2}\Big{|}\Big{\{}i:\tilde{\lambda}_{i}>0\Big{\}}\Big{|}$ quantum link transfers vs. entanglement forging:
@@
IV.2 Modeling qubit movements
@@
  • Variance in the manufacturing process leads to random fluctuations in chip performance, including occasionally defective, unusable qubits (see [4, Figure 1] for a dramatic example). Splitting into chiplets could greatly increase the efficiency of defect rejection. For example, if each qubit has a 1% chance of being defective, the chance of 1000 qubit array emerging with zero defects is roughly $0.4\times 10^{-5}$, likely requiring hundreds of thousands of discarded attempts. As the chance of a 50-qubit array emerging defect-free is about $0.6$, one expects to obtain 20 of these (for 1000 total qubits) in about 34 attempts.
@@
In this Subsection, we combine ideas from the physical channels models described in Subsection III.3 with the network modeling of the previous Subsection IV.1 and the aforementioned chip size vs. noise tradeoffs. We study the noise incurred by a single qubit traversing the graph diameter, calculating the impact of each local SWAP gate or microwave link. We focus on the 27-qubit Falcon processor as chiplet model as illustrated in Figure 2 for several reasons.
@@
    As described in Subsubsection IV.2.2, we have access to and can analyze IBM Falcon processors in detail.
@@
    As concluded in the previous Subsection IV.1, similarity between the chiplet and monolithic architectures motivates analyzing this layout entirely in terms of noise incurred during qubit movements. In this Subsection, we do not directly address bandwidth or latency, which should be only slightly worse than in monolithic heavy-hex.
IV.2.1 A simple model: decaying fidelity
@@
Here we propose a noise model based on standard, simplifying assumptions, in particular that noise is typically depolarizing. As noted in Subsection III.3, the depolarizing noise assumption is not physically realistic for current microwave links, which primarily induce amplitude-damping noise. It is more reasonable for local SWAP gates and compatible with the randomized benchmarking techniques used by IBM to obtain average CX gate errors as reported  [71, 72, 73]. It is not physical realism but a simplifying abstraction that we seek here. In the next Subsubsection, we will address the physical picture in much greater detail.
@@
IV.2.2 Modeling Qubit Movements via Physically Inferred Channels
We seek to address two primary concerns lingering after the previous analyses. First, while analyzing the fidelity of moved qubits gives a broad picture of where distributed quantum computing could show advantages over monolithic, efforts to evaluate specific algorithms or applications might need to consider specifically what kind of errors occur in each setting. In Section IV.4, we use the more detailed models to suggest error detection for links.
Second, we hope to understand whether abstractions are useful rather than hiding important aspects. As in Section IV.1, models based purely on network topology favor as much connectivity as possible. We know however that there are constraints and tradeoff. Section IV.2.1 begins to address these tradeoffs, incorporating them into a model that shows where distributed and monolithic architectures may compete. We still however must acknowledge that the fidelity calculations ignore qualitative differences in the kind of noise from SWAP gates and links. In this Section, we examine the more physically detailed models underlying and abstracted by the fidelity calculations. The results herein validate those abstractions while illuminating their limitations.
@@ IV.2.2 Modeling Qubit Movements via Physically Inferred Channels
    Actual: SWAP chain channel tomography, treating the SWAP-chain as a one-qubit channel as described in Subsection III.3.
@@
    Multiplication: the calculation as in Section IV.2.1.
@@
Since amplitude damping and depolarizing noise do not commute, there is some ambiguity in how one models a quantum process that imposes both. We take the combined channel model corresponding to the continuous limit of both noise types occurring simultaneously. This approach yields a quantum Markov semigroup, a family of channels parameterized continuously by time. The obtained form has a unique and usually full rank fixed point state as long as depolarizing strength is non-zero, a useful property in theoretical analysis and in estimating rates of decay to equilibrium [103, 104]. While the D4 form contains some unitary rotation, because this rotation commutes with the restriction to the fixed point algebra, one may still apply recent results showing exponential decay of relative entropy to the fixed point [105, 106, 107].
@@
  • As expected, the primary contribution to irreversible noise from SWAP gates is depolarizing.
@@
Comparing the D4 model to the simplified model of Equation (10), we find reasonable agreement. We consider the difference in estimated fidelity across values of $m$ and $k$ each ranging from 1 to 20, D4 parameters from each row of Table 1 and a link with $\eta=0.12$. The largest difference between fidelities calculated via D4 and Equation (14) remains below 0.045, which is below 10% of the estimated value from either. Ultimately, the results of this Subsection justify the use of Equation (10) in broadly estimating fidelities. Nonetheless, the more detailed D4 model will likely be useful in fine-grained analysis and designing applications. As discussed in Subsection IV.4, knowing the qualitative form of noise enables simplified and improved error handling.
IV.2.3 Chiplet vs. monolithic performance on qubit movements
In this Section, we compare the fidelity of a qubit that has been moved around a device on monolithic against a chiplet architecture. On September 7, 2021 the 27-qubit ibmq_mumbai reported an average CX fidelity of approximately $F_{CX,Mumbai}=1-0.009$ while the 65-qubit ibmq_brooklyn reported $F_{CX,Brooklyn}=1-0.018$ [108]. This comparison likely represented a much better day for the ibmq_mumbai than for the ibmq_brooklyn and is not a statistically rigorous expectation of typical values. It nonetheless gives a pessimistic of $\Delta_{infid.}$, which we take as an upper bound on how infidelity of monolithic processors may scale.
$\frac{F_{CX,Mumbai}-F_{CX,Brooklyn}}{N_{Brooklyn}-N_{Mumbai}}=\frac{0.991-0.982}{65-27}\approx 0.0002.$ (15)
As a lower bound, we may assume $\Delta_{infid.}=0$, corresponding to no excess noise for larger chips. Since our analysis specifically considers Falcon processors, a class of processors seen online from 2020-2024 [109, 110], we do not attempt to incorporate the broader numerical estimates from [15]. Our bounds on $\Delta_{infid.}=0$ are extremely coarse, representing high uncertainty. Rather than attempt to make detailed predictions about forthcoming technologies, we aim to capture a wide range of possibilities.
@@
We estimate SWAP chain process fidelity using $F_{CX,chip}=1-0.0045$ based on the hybrid surface and Bacon-Shor error threshold [20]. As a note, the infidelity of the hybrid Bacon-Shor threshold is half of the observed average CX infidelity of ibmq_mumbai, $1-F_{CX,Mumbai}$. All other parameters are the same as described in Section IV.2.1. Figure 8 presents SWAP chain process fidelity in a system where the on-chip qubits demonstrate average CX fidelity required for hybrid Bacon-Shor. However, since $F_{link}$ ranges from 0.86-0.96, the described chiplet systems are still not robust enough to implement error correction - either link quality must improve, or even better error rates in the local processor may compensate. Figure 8 includes a comparison of total process fidelity of qubit state transfer on the 27-qubit chiplet architecture vs. monolithic. Fidelity values are noticeably higher than those featured in 7, but the qualitative relationships between ranges chiplet vs. monolithic advantage remain.
@@
IV.3 Modeling fragments of computation
Finally, we depart from graph-driven analyses and consider the fidelity of small computation fragments with or without a link. The network-based comparisons as in Subsection IV.2 avoid this problem by analyzing only the movements of qubits as a building block of computation. That approach may however miss aspects of realistic computation, in which many operations occur simultaneously with qubit movements. Examining realistic computations seems more favorable toward chiplet architectures under the size-noise tradeoffs described in Subsection IV.2, because while the impact of link transfers should remain comparable, the effects of local noise will be amplified when there are many local gates happening alongside qubit movements.

### v4

@@ Abstract
A core challenge for superconducting quantum computers is to scale up the number of qubits in each processor without increasing noise or cross-talk. Distributed quantum computing across small qubit arrays, known as chiplets, can address these challenges in a scalable manner. We propose a chiplet architecture over microwave links with potential to exceed monolithic performance on near-term hardware. Our methods of modeling and evaluating the chiplet architecture bridge the physical and network layers in these processors. We find evidence that distributing computation across chiplets may reduce the overall error rates associated with moving data across the device, despite higher error figures for transfers across links. Preliminary analyses suggest that latency is not substantially impacted, and that at least some applications and architectures may avoid bottlenecks around chiplet boundaries. In the long-term, short-range networks may underlie quantum computers just as local area networks underlie classical datacenters and supercomputers today.
1 Introduction
@@
Quantum computing is at the forefront of emerging technologies, owing to its promise to solve some algorithms dramatically faster than classical counterparts. Quantum computers (QCs) evaluate quantum circuits in a manner similar to classical computers assessing classical circuits. However, quantum information’s ability to leverage superposition, interference, and entanglement is anticipated to provide QCs significant advantage in tasks ranging from security-relevant computing [1] and optimization [2] to simulation of quantum dynamics [3]. To harness the power of quantum computation, quantum bits, or qubits, are needed which are the core individual element for information representation and storage. Useful QCs require many qubits that are interconnected to each other in a low-noise and fault-tolerant manner. Many different architectures have been used to encode qubits, such as superconducting circuits (SCs), trapped ions, neutral atoms, quantum dots, and photons. [4]. To date, SC quantum computers have been one of the most promising platforms [5, 6, 7]. Recent demonstrations of quantum advantage have shown that interference with just over 50 qubits can be harnessed for computational speedups [8, 9, 10]. Additionally, early demonstrations of logical qubits are emerging [11, 12]. However, more and cleaner qubits will be required to exploit quantum computing for practical applications.
@@
State of art QCs, sometimes called Noisy Intermediate Scale Quantum (NISQ) devices, are error prone yet reaching a thousand-qubit capacities [13]. The qubits on these QCs are not error corrected and often have limited connectivity that requires special mapping routines to transform an algorithm into a technology-dependent form [14]. NISQ qubits are characterized by their sensitivity to interactions with their environment and each other, and the gates and measurements required for meaningful computation are implemented with imperfect, error-prone operations. To achieve the realm of usefulness for quantum devices, many more qubits, approximately one million or more, and lower noise figures are essential [15]. Concerningly, as noted in [15], “quantum computer designers face a trade-off between trying to optimize for quantum computers with many physical qubits and quantum computers with very low gate error rate.”
@@
Classical computing saw similar challenges associated with scaling the system-on-chip (SoC) model [27]. Integrated circuit (IC) yield loss is a result of manufacturing defects caused randomly by environmental particulates or by systematic sources such as unavoidable variation or errors during fabrication [28]. Because a larger chip has a higher probability of containing a defect that renders the entire system unusable, fabricating fewer, large SoCs on a wafer results in lower yields than ICs with a smaller wafer footprint. This observation has led to the increase in popularity of chiplet-based architectures for improved cost and flexibility in system design and use [29]. Just as pitfalls were eventually observed in classical computing devices taking a monolithic approach, it is anticipated that continuing to pack qubits on single chip will push the capabilities of manufacturing, cooling, and control technologies to their limit. Thus, to develop QCs containing the millions of qubits required fault-tolerance, we are motivated to pursue a modular approach based on multiple cores working in sync to accomplish computational tasks in a distributed manner.
@@
1.1 Main contributions & where this work fits
A key contribution of this work is to bridge the gap from understanding microwave links at the physical layer to distributed quantum computing applications. Rather than encapsulate network overheads as abstract cost functions, we analyze how noise and costs in both cross-chiplet links and local processing impact the eventual fidelity and feasibility of computations. Our models start and extrapolate from observations of real hardware and experimental physics models. We propose practical, concrete schemes to detect and reduce errors on near-term links and processors. Broad themes in potential applications are evaluated for feasibility on devices with limited connectivity.
@@ 1.1 Main contributions & where this work fits
The structure of this paper is as follows: In Section 2, we review some related literature including analogous approaches to similar questions. Section 3 is about modeling quantum links and local gates at the physical layer, including a comparison of Hamiltonians between optical and microwave links, a review of hardware challenges for local processing, and quantum channel models of relevant operations. We present a quantum channel model that captures realistic noise on both microwave links and local SWAP operations, setting up later comparions. In Section 4, we compare monolithic with chiplet architectures at the network layer, incorporating physical information to derive expected fidelities of qubit transfers in addition to topological metrics. We also derive a detailed channel model that includes both links and local swaps, which we use to propose a simple error detection scheme for the regime of high-quality local processing with lossy links. Finally, we conclude with some open problems and areas for potential follow-up research.
2 Background
@@
2.1 Two major regimes in quantum networking
@@
2.2 Ongoing work in short-range link hardware
@@
In Section 3, we describe the physics and characteristics of these links, comparing them with matter-optical transduction.
2.3 Algorithms and applications
@@
3 Physical components
3.1 On-chip hardware
@@ 3.2 Link hardware
3.3 Modeling links & local SWAP operations
@@
We may alternatively regard a SWAP operation or link as a pair of single-qubit channels, moving a qubit state from the $i$th to $j$th physical qubit. Though this picture is less complete, it vastly simplifies the analysis of multiple moves as in Section 4, characterizing a single-qubit channel via tomography rather than a channel of dimension exponential in the length of the SWAP chain. Empirically, we observed in the experiments for Section 4.2.2 that a long chain model composed from single qubit SWAP tomographies more accurately recovers that of a long chain than does a reduced two-qubit model. We believe this discrepancy reflects cross-talk in nearby measurements or complications in readout error mitigation. The single qubit channel probably characterizes the SWAP more accurately by incorporating less readout noise. Here we again determine an average Choi matrix based on data from the 27-qubit $ibmq\_montreal$:
@@
4 Main Results: modeling distributed architectures & comparing to monolithic
@@
4.1 Connectivities and graph metrics
@@
    As a monolithic comparison point, we consider the standard two-dimensional grid. Each qubit not on the boundary has four neighbors.
@@
    We split the heavy-hex layout into chiplets by replacing some connections with links. This has little effect on topology but may change the error tradeoffs as described in Subsection 4.2. In particular, we partition the chiplet into 80-qubit blocks, each with up to 16 links for a qubit-to-link ratio of at least 10. A smaller version of this layout is shown in Figure 2(c). Unlike the grid case, here we can maintain a reasonable qubit-to-link ratio while keeping the heavy-hex structure across the entire device consistent. We expect topological properties similar to the monolithic heavy-hex but impose a rectangular overall structure, which might be less square than its monolithic counterpart due to the minimum block size.
@@
Figure 4: Comparison of QC topology spectral gap vs. number of qubits for proposed architectures. The spectral gap as in Definition 4.2 quantifies the rate at which random walks on the graph converge toward a uniform distribution on nodes. Graphs with few bottlenecks tend to have higher spectral gaps.
@@
Although topological metrics appear to favor the monolithic grid over other planar layouts, this does not necessarily imply superiority as a quantum computing architecture. As noted in [20], heavy-hex and related layouts achieve lower coordination numbers, taking pressure off of other considerations and potentially enabling techniques for higher fidelity.
@@
4.1.1 Bandwidth considerations for applications
@@
In entanglement forging [61] , an observable on a whole system is rewritten as a linear (not necessarily convex) combination of observables on two subsystems via Schmidt decomposition. Entanglement forging simulates the expectation of an observable $\braket{\mathcal{O}}_{\ket{\psi}}$ on an $n$-qubit system with operator norm error $\epsilon$ via repeated uses of an $n/2$-qubit system. The distributed scheme would use two $n/2$-qubit systems and a number of link uses proportional to the entanglement needed between subsystems. In principle, if each bidirectional link use exchanges up to two qubits of entanglement, the number of link uses is logarithmic in the number of non-zero Schmidt coefficients. For comparability, one should allow a chiplet mapping to approximate the true state within precision $\epsilon$ as well. If we post-select on successful transfers via the two-use code discussed in Section 4.4, then we may convert link uses and noise into a failure probability. The chance that each double-transfer fails is $1/\eta$, where $\eta$ is an exponentiated amplitude damping parameter. In this formulation, we can directly compare the number of retrials expected using $\frac{1}{2}\log_{2}\Big{|}\Big{\{}i:\tilde{\lambda}_{i}>0\Big{\}}\Big{|}$ quantum link transfers vs. entanglement forging:
@@
4.2 Modeling qubit movements
@@
  • Variance in the manufacturing process leads to random fluctuations in chip performance, including occasionally defective, unusable qubits (see [8, Figure 1] for a dramatic example). Splitting into chiplets could greatly increase the efficiency of defect rejection. For example, if each qubit has a 1% chance of being defective, the chance of 1000 qubit array emerging with zero defects is roughly $0.4\times 10^{-5}$, likely requiring hundreds of thousands of discarded attempts. As the chance of a 50-qubit array emerging defect-free is about $0.6$, one expects to obtain 20 of these (for 1000 total qubits) in about 34 attempts.
@@
In this Subsection, we combine ideas from the physical channels models described in Subsection 3.3 with the network modeling of the previous Subsection 4.1 and the aforementioned chip size vs. noise tradeoffs. We study the noise incurred by a single qubit traversing the graph diameter, calculating the impact of each local SWAP gate or microwave link. We focus on the 27-qubit Falcon processor as chiplet model as illustrated in Figure 2 for several reasons.
@@
    As described in Subsubsection 4.2.2, we have access to and can analyze IBM Falcon processors in detail.
@@
    As concluded in the previous Subsection 4.1, similarity between the chiplet and monolithic architectures motivates analyzing this layout entirely in terms of noise incurred during qubit movements. In this Subsection, we do not directly address bandwidth or latency, which should be only slightly worse than in monolithic heavy-hex.
4.2.1 A simple model: decaying fidelity
@@
Here we propose a noise model based on standard, simplifying assumptions, in particular that noise is typically depolarizing. As noted in Subsection 3.3, the depolarizing noise assumption is not physically realistic for current microwave links, which primarily induce amplitude-damping noise. It is more reasonable for local SWAP gates and compatible with the randomized benchmarking techniques used by IBM to obtain average CX gate errors as reported  [71, 72, 73]. It is not physical realism but a simplifying abstraction that we seek here. In the next Subsubsection, we will address the physical picture in much greater detail.
@@
4.2.2 Modeling qubit movements via physically inferred channels
We seek to address two primary concerns lingering after the previous analyses. First, while analyzing the fidelity of moved qubits gives a broad picture of where distributed quantum computing could show advantages over monolithic, efforts to evaluate specific algorithms or applications might need to consider specifically what kind of errors occur in each setting. In Section 4.4, we use the more detailed models to suggest error detection for links.
Second, we hope to understand whether abstractions are useful rather than hiding important aspects. As in Section 4.1, models based purely on network topology favor as much connectivity as possible. We know however that there are constraints and tradeoff. Section 4.2.1 begins to address these tradeoffs, incorporating them into a model that shows where distributed and monolithic architectures may compete. We still however must acknowledge that the fidelity calculations ignore qualitative differences in the kind of noise from SWAP gates and links. In this Section, we examine the more physically detailed models underlying and abstracted by the fidelity calculations. The results herein validate those abstractions while illuminating their limitations.
@@ 4.2.2 Modeling qubit movements via physically inferred channels
    Actual: SWAP chain channel tomography, treating the SWAP-chain as a one-qubit channel as described in Subsection 3.3.
@@
    Multiplication: calculated as in Section 4.2.1.
@@
Since amplitude damping and depolarizing noise do not commute, there is some ambiguity in how one models a quantum process that imposes both. We take the combined channel model corresponding to the continuous limit of both noise types occurring simultaneously. This approach yields a quantum Markov semigroup, a family of channels parameterized continuously by time.
@@
  • As expected, the primary contribution to irreversible SWAP gate noise is depolarizing.
@@
Comparing the D4 model to the simplified model of Equation (10), we find reasonable agreement. We consider the difference in estimated fidelity across values of $m$ and $k$ each ranging from 1 to 20, D4 parameters from each row of Table 1 and a link with $\eta=0.12$. The largest difference between fidelities calculated via D4 and Equation (14) remains below 0.045, which is below 10% of the estimated value from either. Ultimately, the results of this Subsection justify the use of Equation (10) in broadly estimating fidelities. Nonetheless, the more detailed D4 model will likely be useful in fine-grained analysis and designing applications. As discussed in Subsection 4.4, knowing the qualitative form of noise enables simplified and improved error handling.
4.2.3 Chiplet vs. monolithic performance on qubit movements
In this Section, we compare the fidelity of a qubit that has been moved around a device on monolithic against a chiplet architecture. On September 7, 2021 the 27-qubit ibmq_mumbai reported an average CX fidelity of approximately $F_{CX,Mumbai}=1-0.009$ while the 65-qubit ibmq_brooklyn reported $F_{CX,Brooklyn}=1-0.018$ [103]. This comparison likely represented a much better day for the ibmq_mumbai than for the ibmq_brooklyn and is not a statistically rigorous expectation of typical values. It nonetheless gives a pessimistic estimate of $\Delta_{infid.}$, which we take as an upper bound on how infidelity of monolithic processors may scale.
$\begin{split}\frac{F_{CX,Mumbai}-F_{CX,Brooklyn}}{N_{Brooklyn}-N_{Mumbai}}&=\frac{0.991-0.982}{65-27}\\ &\approx 0.0002.\end{split}$ (15)
As a lower bound, we may assume $\Delta_{infid.}=0$, corresponding to no excess noise for larger chips. Since our analysis specifically considers Falcon processors, a class of processors seen online from 2020-2024 [104, 105], we do not attempt to incorporate the broader numerical estimates from [15]. Our bounds on $\Delta_{infid.}$ are coarse, representing high uncertainty. Rather than attempt to make detailed predictions about forthcoming technologies, we aim to capture a wide range of possibilities.
@@
We estimate SWAP chain process fidelity using $F_{CX,chip}=1-0.0045$ based on the hybrid surface and Bacon-Shor error threshold [20]. As a note, the infidelity of the hybrid Bacon-Shor threshold is half of the observed average CX infidelity of ibmq_mumbai, $1-F_{CX,Mumbai}$. All other parameters are the same as described in Section 4.2.1. Figure 8 presents SWAP chain process fidelity in a system where the on-chip qubits demonstrate average CX fidelity required for hybrid Bacon-Shor. However, since $F_{link}$ ranges from 0.86-0.96, the described chiplet systems are still not robust enough to implement error correction - either link quality must improve, or even better error rates in the local processor may compensate. Figure 8 includes a comparison of total process fidelity of qubit state transfer on the 27-qubit chiplet architecture vs. monolithic. Fidelity values are noticeably higher than those featured in 7, but the qualitative relationships between ranges chiplet vs. monolithic advantage remain.
@@
4.3 Modeling fragments of computation
Finally, we depart from graph-driven analyses and consider the fidelity of small computation fragments with or without a link. The network-based comparisons as in Subsection 4.2 avoid this problem by analyzing only the movements of qubits as a building block of computation. That approach may however miss aspects of realistic computation, in which many operations occur simultaneously with qubit movements. Examining realistic computations seems more favorable toward chiplet architectures under the size-noise tradeoffs described in Subsection 4.2, because while the impact of link transfers should remain comparable, the effects of local noise will be amplified when there are many local gates happening alongside qubit movements.